The balanced combustion reaction for C6H6 is

2 C6H6(l) + 15 O2 (g) -----> 12 CO2 (g) + 6 H2O(l) + 6542 kJ

If 8.800 g C6H6 is burned and the heat produced from the burning is added to 5691 g of water at 21 C, What is the final temperature of the water?

final temperature= ___________ C
Answer #1

The balanced equation is

2 C6H6(l) + 15 O2(g) -------------- 12 CO2(g) + 6 H2O + 6542 KJ

mass of C6H6 = 8.800 g

molar mass of C6H6 = 78.11 g/mole

number of moles of C6H6 = mass / molar mass = 8.800 / 78.11 = 0.1127 mole

number of moles of C6H6 = 0.1127 mole

according to equation

heat generated from 2 moles of C6H6 = 6542 KJ

heat generated from 0.1127 moles of C6H6 = ?

                                                                      = 0.1127 x 6542/2 = 368.64 KJ

Heat generated = 368.64 KJ

q= 368.64 KJ = 368640 J

mass of water = m= 5691 g

specific heat capacity of water = c= 4.184 J/gk

Initial temperature = T1= 21C

Final temperature= T2 = ?

q= ms( T2-T1)

368640 = 5691 x 4.184 x ( T2 - 21)

T2 - 21= 15.48

T2= 36.48 C

Final temperature = T2 = 36.48 C